OpenStudy (anonymous):

what is the probability of rolling a prime number or a perfect square (addition) with a standard pair of dice

OpenStudy (anonymous):

hmmm so you are rolling one dice or 2 dies at once?

OpenStudy (anonymous):

2 dice

OpenStudy (anonymous):

K so if we are throwing 2 dies the minimum sum is 2 and the maximum sum is 12 So what numbers btwn 2 and 12 including 2 and 12 are prime number?

OpenStudy (anonymous):

im not sure what a prime number is

OpenStudy (anonymous):

prime number is a number that doesnt have any factors except 1 and itself For example 3 - it can only be divided by 1 and 3 Another example is 5 - 5 doesnt have any divisors except 5 and 1

OpenStudy (anonymous):

3, 5, 7, 11,

OpenStudy (anonymous):

correct ... also less noticeable 2 2 can only be divided by 1 and 2

OpenStudy (anonymous):

So the prime numbers are 2,3,5,7,11

OpenStudy (anonymous):

A perfect square is a number that can be squared and will result in an integer or another way of looking at it is is a number that has 2 of the same factors so for example 25=5*5 16=4*4 Now basically we can determine our perfect squares as follows 1*1=1 2*2=4 3*3=9 4*4=16 etc So which number in btwn 2 and 12 are perfect squares?

OpenStudy (anonymous):

1 2 and 3

OpenStudy (anonymous):

1 ,2 and 3 are the factors 1,4,9 and 16 are known as perfect squares

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

So 4 and 9 are perfect squares in btwn 2 and 12

OpenStudy (anonymous):

Now lets draw a chart of all the different options

OpenStudy (anonymous):